OpenStudy (anonymous):

Please help! If a varies jointly as b and c, and a=27 when b=18 and c=3, find a when b=6 and c=1/3.

OpenStudy (anonymous):

\[\frac{27}a{}=\frac{18*3}{6*(\frac{1}{3})}\]

OpenStudy (anonymous):

can u solve now?

OpenStudy (anonymous):

I think so. Thank you for the help!

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

54/2=27? Now what?

OpenStudy (anonymous):

a=1

OpenStudy (anonymous):

27a=27 a=27/27 a=1

OpenStudy (anonymous):

Ohhh. I get it. Divide by 2. Thank you!

OpenStudy (anonymous):

ty